# The strength of a geometric simplex is a key ingredient in a polynomial-time classification of unordered point clouds by Lipschitz continuous invariants

### v1

@@ 3 Proofs of strength properties in all dimensions
Since each $(n+1)\times(n+1)$ submatrix includes one entry $d_{ij}^{2}$, we exclude all products with this entry, which are multiplied by the removed $d_{ij}^{2}$ from $\hat{D}$ and hence were counted in $\alpha d_{ij}^{4}$. Hence we can replace $d_{02}=d_{02}$ with 0 and get $\beta=\left|\begin{array}[]{cccc}0&1&1&1\\ 1&d_{10}^{2}&0&d_{13}^{2}\\ 1&0&d_{21}^{2}&d_{23}^{2}\\ 1&d_{30}^{2}&d_{32}^{2}&0\end{array}\right|+\left|\begin{array}[]{cccc}0&1&1&1\\ 1&d_{01}^{2}&0&d_{03}^{2}\\ 1&0&d_{12}^{2}&d_{13}^{2}\\ 1&d_{31}^{2}&d_{32}^{2}&0\end{array}\right|$. Up to a permutation of indices, each submatrix can be rewritten with the diagonal that has one original $d_{ij}^{2}$ (now replaced with 0), while all other diagonal elements are initial zeros. The example above gives $\beta=-\left|\begin{array}[]{cccc}0&1&1&1\\ 1&0&d_{21}^{2}&d_{23}^{2}\\ 1&d_{10}^{2}&0&d_{13}^{2}\\ 1&d_{30}^{2}&d_{32}^{2}&0\end{array}\right|-\left|\begin{array}[]{cccc}0&1&1&1\\ 1&0&d_{12}^{2}&d_{13}^{2}\\ 1&d_{01}^{2}&0&d_{03}^{2}\\ 1&d_{31}^{2}&d_{32}^{2}&0\end{array}\right|$. Similarly to the argument for the determinant $\alpha$, the sum $\beta d_{ij}$ contains $2r_{n+1}$ products of the total degree 2n12n-1, so $\dfrac{|\beta|d_{ij}}{p^{2n-1}(A)}\leq 2r_{n+1}\left(\dfrac{2}{n}\right)^{2n-1}$ by Lemma 3.9. Then the required inequality follows: $\left|\dfrac{\partial\det\hat{D}}{\partial d_{ij}}\right|\dfrac{1}{p^{2n-1}(A)}=\dfrac{|4\alpha d_{ij}^{3}+2\beta d_{ij}|}{p^{2n-1}(A)}\leq(4r_{n}+4r_{n+1})\left(\dfrac{2}{n}\right)^{2n-1}$. ∎
@@
For $n=3$, the squared volume is $\mathrm{vol}^{2}(A)=\dfrac{1}{288}\left|\begin{array}[]{ccccc}0&1&1&1&1\\ 1&0&d_{01}^{2}&d_{02}^{2}&d_{03}^{2}\\ 1&d_{10}^{2}&0&d_{12}^{2}&d_{13}^{2}\\ 1&d_{20}^{2}&d_{21}^{2}&0&d_{23}^{2}\\ 1&d_{30}^{2}&d_{31}^{2}&d_{32}^{2}&0\end{array}\right|$.
@@
Similarly to the case $n=2$, the mean value theorem [5] for the strength $\sigma(A)=\dfrac{\mathrm{vol}^{2}(A)}{p^{2n-1}(A)}$ implies that $|\sigma(A)-\sigma(B)|\leq 2\varepsilon\sup|\nabla\sigma|\leq 2\varepsilon\sqrt{\sum\limits_{i\neq j}\sup\left|\dfrac{\partial\sigma}{\partial d_{ij}}\right|^{2}}\leq 2\varepsilon\sqrt{\dfrac{n(n+1)}{2}}\max\limits_{i\neq j}\sup\left|\dfrac{\partial\sigma}{\partial d_{ij}}\right|.$ To find an upper bound of $\left|\dfrac{\partial\sigma}{\partial d_{ij}}\right|$, we initially ignore the numerical factor in the square volume $\mathrm{vol}^{2}(A)=\dfrac{(-1)^{n-1}}{2^{n}(n!)^{2}}\det\hat{D}$ and differentiate $\det\hat{D}\cdot\dfrac{1}{p^{2n-1}(A)}$ by the product rule: $\dfrac{\partial}{\partial d_{ij}}\left(\dfrac{\det\hat{D}}{p^{2n-1}(A)}\right)=\dfrac{\partial\det\hat{D}}{\partial d_{ij}}\cdot\dfrac{1}{p^{2n-1}(A)}-\dfrac{\det\hat{D}}{p^{2n}(A)}\cdot\dfrac{2n-1}{2}$. Lemmas 3.11 and 3.12 below imply the upper bound $\left|\dfrac{\partial}{\partial d_{ij}}\left(\dfrac{\det\hat{D}}{p^{2n-1}(A)}\right)\right|\leq(4r_{n}+4r_{n+1})\left(\dfrac{2}{n}\right)^{2n-1}+\left(n-\dfrac{1}{2}\right)r_{n+2}\left(\dfrac{2}{n}\right)^{2n}<\left(4r_{n}+4r_{n+1}+2r_{n+2}\right)\left(\dfrac{2}{n}\right)^{2n-1}$. Taking into account the factors $\dfrac{(-1)^{n-1}}{2^{n}(n!)^{2}}$ in $\mathrm{vol}^{2}(A)$ and $\sqrt{\dfrac{n(n+1)}{2}}$ for estimating the length of the gradient $\nabla\sigma$ of $\dfrac{n(n+1)}{2}$ first order partial derivatives $\dfrac{\partial\sigma}{\partial d_{ij}}$, we get $|\sigma(A)-\sigma(B)|\leq 2\varepsilon\sqrt{\dfrac{n(n+1)}{2}}\max\limits_{i\neq j}\sup\left|\dfrac{\partial\sigma}{\partial d_{ij}}\right|\leq 2\varepsilon c_{n}$ for the upper bound $c_{n}=\dfrac{\sqrt{n(n+1)/2}}{2^{n}(n!)^{2}}\left(4r_{n}+4r_{n+1}+2r_{n+2}\right)\left(\dfrac{2}{n}\right)^{2n-1}=(2r_{n}+2r_{n+1}+r_{n+2})\dfrac{2^{n-0.5}\sqrt{n+1}}{(n!)^{2}n^{2n-1.5}}$. The estimate $r_{n}\leq\dfrac{n!}{2}$ gives $c_{n}<\dfrac{n!}{2}(2+2(n+1)+(n+1)(n+2))\dfrac{2^{n-0.5}\sqrt{n+1}}{(n!)^{2}n^{2n-1.5}}=(n^{2}+5n+6)\sqrt{n+1}\dfrac{2^{n-1.5}}{(n!)n^{2n-1.5}}=\left(1+\dfrac{5}{n}+\dfrac{6}{n^{2}}\right)\sqrt{1+\dfrac{1}{n}}\dfrac{2^{n-1.5}}{(n!)n^{2n-4}}$. For $n\geq 3$, we get $1+\dfrac{5}{n}+\dfrac{6}{n^{2}}\leq\dfrac{10}{3}$ and $\sqrt{1+\dfrac{1}{n}}\leq\dfrac{2}{\sqrt{3}}$, so $c_{n}<\dfrac{20}{3\sqrt{3}}\cdot\dfrac{2^{n-1.5}}{(n!)n^{2n-4}}=\dfrac{5\sqrt{2}}{3\sqrt{3}}\dfrac{2^{n}}{(n!)n^{2n-4}}<\dfrac{2^{n+0.5}}{(n!)n^{2n-4}}=b_{n}$. ∎

### v2

@@ 3 Proofs of strength properties in all dimensions
Since each $(n+1)\times(n+1)$ submatrix includes one entry $d_{ij}^{2}$, we exclude all products with this entry, which are multiplied by the removed $d_{ij}^{2}$ from $\hat{D}$ and hence were counted in $\alpha d_{ij}^{4}$. Hence we can replace $d_{02}=d_{02}$ with 0 and get $\beta=\left|\begin{array}[]{cccc}0&1&1&1\\ 1&d_{10}^{2}&0&d_{13}^{2}\\ 1&0&d_{21}^{2}&d_{23}^{2}\\ 1&d_{30}^{2}&d_{32}^{2}&0\end{array}\right|+\left|\begin{array}[]{cccc}0&1&1&1\\ 1&d_{01}^{2}&0&d_{03}^{2}\\ 1&0&d_{12}^{2}&d_{13}^{2}\\ 1&d_{31}^{2}&d_{32}^{2}&0\end{array}\right|$. Up to a permutation of indices, each submatrix can be rewritten with the diagonal that has one original $d_{ij}^{2}$ (now replaced with 0), while all other diagonal elements are initial zeros. The example above gives $\beta=-\left|\begin{array}[]{cccc}0&1&1&1\\ 1&0&d_{21}^{2}&d_{23}^{2}\\ 1&d_{10}^{2}&0&d_{13}^{2}\\ 1&d_{30}^{2}&d_{32}^{2}&0\end{array}\right|-\left|\begin{array}[]{cccc}0&1&1&1\\ 1&0&d_{12}^{2}&d_{13}^{2}\\ 1&d_{01}^{2}&0&d_{03}^{2}\\ 1&d_{31}^{2}&d_{32}^{2}&0\end{array}\right|$. Similarly to the argument for the determinant $\alpha$, the sum $\beta d_{ij}$ contains $2r_{n+1}$ products of the total degree 2n12n-1, so $\dfrac{|\beta|d_{ij}}{p^{2n-1}(T)}\leq 2r_{n+1}\left(\dfrac{2}{n}\right)^{2n-1}$ by Lemma 3.2. Then the required inequality follows: $\left|\dfrac{\partial\det\hat{D}}{\partial d_{ij}}\right|\dfrac{1}{p^{2n-1}(T)}=\dfrac{|4\alpha d_{ij}^{3}+2\beta d_{ij}|}{p^{2n-1}(T)}\leq(4r_{n}+4r_{n+1})\left(\dfrac{2}{n}\right)^{2n-1}$. ∎
@@
For $n=3$, the squared volume is $\mathrm{vol}^{2}(T)=\dfrac{1}{288}\left|\begin{array}[]{ccccc}0&1&1&1&1\\ 1&0&d_{01}^{2}&d_{02}^{2}&d_{03}^{2}\\ 1&d_{10}^{2}&0&d_{12}^{2}&d_{13}^{2}\\ 1&d_{20}^{2}&d_{21}^{2}&0&d_{23}^{2}\\ 1&d_{30}^{2}&d_{31}^{2}&d_{32}^{2}&0\end{array}\right|$. Similarly to the case $n=2$, the mean value theorem [elliott2012probabilistic] for the strength $\sigma(T)=\dfrac{\mathrm{vol}^{2}(T)}{p^{2n-1}(T)}$ implies that $|\sigma(T)-\sigma(Q)|\leq 2\varepsilon\sup|\nabla\sigma|\leq 2\varepsilon\sqrt{\sum\limits_{i\neq j}\sup\left|\dfrac{\partial\sigma}{\partial d_{ij}}\right|^{2}}\leq 2\varepsilon\sqrt{\dfrac{n(n+1)}{2}}\max\limits_{i\neq j}\sup\left|\dfrac{\partial\sigma}{\partial d_{ij}}\right|.$ To find an upper bound of $\left|\dfrac{\partial\sigma}{\partial d_{ij}}\right|$, we initially ignore the numerical factor in the square volume $\mathrm{vol}^{2}(T)=\dfrac{(-1)^{n-1}}{2^{n}(n!)^{2}}\det\hat{D}$ and differentiate $\det\hat{D}\cdot\dfrac{1}{p^{2n-1}(T)}$ by the product rule:
$\dfrac{\partial}{\partial d_{ij}}\left(\dfrac{\det\hat{D}}{p^{2n-1}(T)}\right)=\dfrac{\partial\det\hat{D}}{\partial d_{ij}}\cdot\dfrac{1}{p^{2n-1}(T)}-\dfrac{\det\hat{D}}{p^{2n}(T)}\cdot\dfrac{2n-1}{2}.$
Lemmas 3.3 and 3.4 imply the upper bound
$\left|\dfrac{\partial}{\partial d_{ij}}\left(\dfrac{\det\hat{D}}{p^{2n-1}(T)}\right)\right|\leq(4r_{n}+4r_{n+1})\left(\dfrac{2}{n}\right)^{2n-1}+\left(n-\dfrac{1}{2}\right)r_{n+2}\left(\dfrac{2}{n}\right)^{2n}<$
$\left(4r_{n}+4r_{n+1}+2r_{n+2}\right)\left(\dfrac{2}{n}\right)^{2n-1}$. Taking into account the factors $\dfrac{(-1)^{n-1}}{2^{n}(n!)^{2}}$ in $\mathrm{vol}^{2}(T)$ and $\sqrt{\dfrac{n(n+1)}{2}}$ for estimating the length of the gradient $\nabla\sigma$ of $\dfrac{n(n+1)}{2}$ first order partial derivatives $\dfrac{\partial\sigma}{\partial d_{ij}}$, we get
$|\sigma(T)-\sigma(Q)|\leq 2\varepsilon\sqrt{\dfrac{n(n+1)}{2}}\max\limits_{i\neq j}\sup\left|\dfrac{\partial\sigma}{\partial d_{ij}}\right|\leq 2\varepsilon c_{n}$
for the upper bound
$\begin{array}[]{l}c_{n}=\dfrac{\sqrt{n(n+1)/2}}{2^{n}(n!)^{2}}\left(4r_{n}+4r_{n+1}+2r_{n+2}\right)\left(\dfrac{2}{n}\right)^{2n-1}=\\ =(2r_{n}+2r_{n+1}+r_{n+2})\dfrac{2^{n-0.5}\sqrt{n+1}}{(n!)^{2}n^{2n-1.5}}.\end{array}$
The estimate $r_{n}\leq\dfrac{n!}{2}$ gives
$\begin{array}[]{l}c_{n}<\dfrac{n!}{2}\Big(2+2(n+1)+(n+1)(n+2)\Big)\dfrac{2^{n-0.5}\sqrt{n+1}}{(n!)^{2}n^{2n-1.5}}=\\ =(n^{2}+5n+6)\sqrt{n+1}\dfrac{2^{n-1.5}}{(n!)n^{2n-1.5}}=\left(1+\dfrac{5}{n}+\dfrac{6}{n^{2}}\right)\sqrt{1+\dfrac{1}{n}}\dfrac{2^{n-1.5}}{(n!)n^{2n-4}}.\end{array}$
For $n\geq 3$, we get $1+\dfrac{5}{n}+\dfrac{6}{n^{2}}\leq\dfrac{10}{3}$ and $\sqrt{1+\dfrac{1}{n}}\leq\dfrac{2}{\sqrt{3}}$, so
$c_{n}<\dfrac{20}{3\sqrt{3}}\cdot\dfrac{2^{n-1.5}}{(n!)n^{2n-4}}=\dfrac{5\sqrt{2}}{3\sqrt{3}}\dfrac{2^{n}}{(n!)n^{2n-4}}<\dfrac{2^{n+0.5}}{(n!)n^{2n-4}}=b_{n}.$
Proof of Lipschitz constants for the signed strength in Theorem 2.4(b).
If simplices $T,Q$ have $\mathrm{sign}(T)=\mathrm{sign}(Q)$, then $|s(T)-s(Q)|=|\sigma(T)-\sigma(Q)|\leq 2\lambda_{n}\varepsilon$.
If simplices $T,Q$ have opposite signs, we will prove that $|s(T)-s(Q)|\leq 4\lambda_{n}\varepsilon$. Any simplex $T$ can be connected to its $\varepsilon$-perturbation $Q$ by a straight-line deformation through simplices $Q_{t}\subset\mathbb{R}^{n}$, where $t\in[0,1]$, $Q_{0}=T$, and $Q_{1}=Q$.
There is an intermediate value $t\in(0,1)$ such that $\sigma(Q_{t})=0=\mathrm{sign}(Q_{t})$. Then both $T,Q$ are deformations of the intermediate simplex $Q_{t}$ such that all corresponding vertices are perturbed up to Euclidean distances $t\varepsilon$ and $(1-t)\varepsilon$, respectively. The inequalities $|s(T)|=|\sigma(T)-\sigma(Q_{t})|\leq 2\lambda_{n}t\varepsilon$ and $|s(Q)|=|\sigma(Q_{t})-\sigma(Q)|\leq 2\lambda_{n}(1-t)\varepsilon$ imply that $|s(T)-s(Q)|\leq|\sigma(T)-\sigma(Q_{t})|+|\sigma(Q_{t})-\sigma(Q)|\leq 2\lambda_{n}\varepsilon$. ∎
In conclusion, the strength of a simplex provides a Lipschitz continuous analog of a volume, which is measured in the units of original point coordinates. We thank Janos Pach for helpful discussions of this paper leading to forthcoming joint work.
Data availability statement. All data for this research is included in the paper.